You lower a 3.10-kg box to the floor from a shelf 1.12m above the floor. What is the magnitude of the change in potential energy of the box? Does it increase or decrease?
Answer #1

Potential energy = mass x acceleration due to gravity x height above floor

= mgh

Here, final potential energy = mg(0) = 0 J

Initial potential energy = mg(1.12) = (3.10)(9.81)(1.12) = 34.06 J

Change in potential energy = Final potential energy - initial potential energy = 0 - 34.06 = - 34.06 J

Since change is negative, potential energy decreases